Volunteer's View - late September to early October

Recent reports from Darrell Marsh and Graham Clarke…

September 19th: A working party at BH with me, Bob, Martin, Nigel, John J, John G, Mike and Howard attending, some for part of the day and some all day.

Mike had a refresher on both forklifts and Nigel and Howard were trained on the Doosan. All passed.

Other work carried out today included the following: John J and Bob worked on the lagging for the collector drum on PU417. Hooks were sourced and fitted to two pieces. These were then fitted to the back piece on the drum. Further hooks will need to be removed from old lagging and more copper wire bought. Work on this will continue at the next working party on Wednesday 2nd October. Martin separated and cleaned the last three PU feet. New pads can now be fitted. New rubber is still to be sourced. As the weather was good Nigel continued with the cab doors for the cab outside as there was more sanding to be done. John G made a welcome return to the midweek working parties and spent the day painting cab components for 55008 and did most of the tea making. Along with supervising the volunteers I made a start on the yearly stocktake.

October 2nd: The latest midweek working day with me, Bob, Mark, Nigel, Trevor, John J, and Geoff attending. Overall I think we had a good day though it would have been useful if I could have cloned two more of me. During the day the following was carried out:

Bob and John carried on with the lagging on the collector drum on PU417. Mark tidied stores floor to allow fork truck in to retrieve radiators, assisted me loading van for SVR and worked on some components for the cab. Nigel continues with cab and nose end doors on 55008′s cab Trevor search for bolts etc in readiness to start fitting the radiators and removed the two broken gutter drain pipes on D9015. Geoff degreased the buffer taken apart by Dave on Saturday. I loaded van with sales stock for the SVR gala, helped Trevor finding bolts, met Greg and David to sort website and met a supplier who can supply a rubber gasket we need for the PU feet the midweek team have been working on.

Darrell Marsh

Now a report from Graham Clarke:

From my own point of view, I have made further progress with connecting, continuity testing and insulation resistance testing the end to end wiring for the traction motor blowers and vacuum exhauster wiring. Everything is satisfactory. I’ve also started getting ready to connect up the fuel pumps. I need some metric to imperial conduit adapters to go into the motor terminal boxes. The new conduit adapter is metric and larger in diameter than the old imperial thread in the terminal box. Both Nigels and Beth worked on 55015’s lower green band, cant rail and doors for 55008.

Last Saturday, Mike and I got the radiators down ready to fit into 55015, re-stacked the spare oil radiators and found enough nose compartment air filters for 55015. New ones would cost about £2,500, so it was a profitable afternoon! They need some handles removing to make them fit, which is quite easy. After Mike left, I checked the traction motor cabling, because I thought I had wired some of the connection boxes under 55015 back to front, but I haven’t. Phew!

Lastly, the new traction current ammeters have arrived for 55015, courtesy of Mike and Trent Instruments.

A very big thank you to everyone who contributed over the last couple of weeks.